In an ac circuit it is containing resistance of 15 ohm and inductance of 0.3H in series with 50 Hz ac supply voltage.If the voltage across resistance and inductance is 10 volt and 15 volt respectively.Then determine the total input voltage of the circuit

1
Expert's answer
2022-03-09T10:32:31-0500

The vector diagram of the resultant voltage for AC RL series circuit is shown below


The total input voltage

"V=\sqrt{V_R^2+V_L^2}=\sqrt{10^2+15^2}=18\:\rm V"